OpenStudy (anonymous):

in a tiriangle abc angle b is 5 times angle a and angle c is 2 degress less tahn 4 times angle a find the size of the angles

OpenStudy (anonymous):

all the angles in a triangle add up to 180. write an equation of all angles in terms of angle a and solve

OpenStudy (anonymous):

i am not sure how to write an equation

OpenStudy (anonymous):

b = 5a c=4a-2 thus a + 5a+ (4a - 2) = 180 a+5a+4a-2=180 10a-2=180 10a=182 a=18.2 Since we have a, you can find the other angles

OpenStudy (anonymous):

thanks i got them

OpenStudy (anonymous):

"all the angles in a triangle add up to 180." Its correct only when the triangle is in Euclidian Space. Remember that. Its very important thing.